# What is the Relationship between Tensor Factorizations and Circuits (and How Can We Exploit it)?

### v1

@@ 2.2 Hierarchical Tensor Factorizations are Deep Circuits
Opportunity 1. A wider choice of factorization structures Refer to caption Figure 5: Region nodes can be shared between partitionings in a DAG RG. Def. 4 allows for arbitrary DAGs and arbitrarily scoped-regions, while hierarchical tensor factorizations are usually presented in terms of RGs with a tree structure having region leaves containing exactly one variable (e.g., the RG in Fig. 4), which are sometimes called dimension trees or mode cluster trees (Grasedyck, 2010) in the tensor factorization community, and often vtree in circuit literature (Pipatsrisawat & Darwiche, 2008; Kisa et al., 2014; Wedenig et al., 2024a). More intricate RG structures can increase the expressiveness as well as flexibility in building deep circuits/hierarchical factorizations. Intuitively, we can share factor matrices among multiple factorizations, and therefore reduce the number of model parameters, making a more space-efficient implementation possible. See Peharz et al. (2020c; a) for more details. We provide an example of such a RG in Fig. 9 and Fig. 5. The reader can check that this RG encodes the hierarchical scope partitioning of the decomposable circuit in figure Fig. 1. Fig. 8 then illustrates a fragment of this RG and shows how tensor factorizations conforming to fragments of this RG can be constructed as circuits in Section 4. The circuit literature provides several ways to build RGs that are suitable for certain data modalities (e.g., image, sequence and tabular data), which can also be adaptively learned from data. Section 4.1 provides an overview of such techniques.
@@
Note that instantiating tensor factorizations from RGs defined as above preserve decomposability, and that circuits built from RGs in the literature are typically also smooth (Def. 3). Hierarchical Tucker and its variants are also smooth and decomposable and therefore support the tractable computation of a number of (probabilistic) inference tasks (Section 3). These hierarchical factorizations (and the corresponding deep circuits) that follow a tree-shaped RG with univariate leaves satisfy an additional structural property, called structured-decomposability. Structured decomposability enables the tractable computation of harder operations for which smoothness and decomposability are not enough. For instance, squaring particular tensor factorizations formalized in the graphical language of tensor networks, known as the Born rule in physics (Feynman, 1987; Glasser et al., 2019) (see also Section 2.4). We define structured decomposability below.
@@
We can check that hierarchical Tucker yields a structured decomposable circuit easily, as it is obtained by stacking Tucker factorizations (which are computed by decomposable circuits) based on a tree RG, which in turn synchronizes all product units to decompose in the same way. We emphasize that eliciting the few structural properties that can explain the tractable computation of many different quantities of interest can help save effort aimed at (re)discovering and (re)engineering algorithms for specific hierarchical factorizations.
Opportunity 2. Efficient Compositional Operations over Factorizations Given one or more tensor factorizations appearing as operands in a computation of interest, how can we automatically devise a tractable algorithm for it without having to materialize the exponentially large tensor operands? The circuit literature holds the answer and offers other structural properties that can unlock the tractable computation of many complex inference scenarios, in a reusable fashion. E.g., when two deep circuits conform to the same tree RG, they are said to be compatible (Vergari et al., 2021). Given two hierarchical tensor factorizations $p$ and $q$ over $\bm{\mathrm{X}}$, if they are compatible one can compute general expectations of the form $\sum\nolimits_{\bm{\mathrm{x}}}p(\bm{\mathrm{x}})q(\bm{\mathrm{x}}),$ (expectations) in closed form in time $\mathcal{O}(|p||q|)$, where $|p|$ and $|q|$ are the size of the corresponding circuits encoding such factorizations.
@@
On the other hand, maximization problems as in maximum-a-posteriori inference $\max\nolimits_{\bm{\mathrm{y}}}p(\bm{\mathrm{y}},\bm{\mathrm{E}}=\bm{\mathrm{e}}),$ (MAP inference) where $\bm{\mathrm{e}}$ is the evidence assignment to variables $\bm{\mathrm{E}}\subset\bm{\mathrm{X}}$, and $\bm{\mathrm{y}}$ is the assignment to the remaining variables $\bm{\mathrm{Y}}=\bm{\mathrm{X}}\setminus\bm{\mathrm{E}}$ for which we want to maximize $p$, can be solved exactly and efficiently if $p$ is a decomposable circuit that supports an additional property, determinism (Darwiche, 2009). In a nutshell, sum units in a deterministic circuit receive inputs from functions with disjoint support (see Choi et al. (2020) for details). While determinism is a consolidated property in the circuit literature, it is off the radar for (hierarchical) tensor factorizations. Furthermore, the circuit literature provides a systematic way to quickly devise the tractability conditions for a given mathematical expression that involves sums, products, powers, exponentials and logarithms, and therefore automatically distill corresponding tractable algorithms (Vergari et al., 2021). For example, if one wants to compute the Rényi’s $\alpha$-divergence between two factorizations $p$ and $q$ over variables $\bm{\mathrm{X}}$, for $\alpha\in\mathbb{N}$, defined as $(1-\alpha)^{-1}\log\sum\nolimits_{\bm{\mathrm{x}}}p^{\alpha}(\bm{\mathrm{x}})q^{1-\alpha}(\bm{\mathrm{x}}),\$ ($\alpha$-divergence) then this can be done quickly if $p$ and $q$ can be represented as smooth, decomposable and compatible circuits and $q$ is also deterministic. Vergari et al. (2021) shows how to automatically distill the tractable computation of more information-theoretic quantities.
@@ 2.3 Representing Circuits in a Tensorized Formalism
Representing (hierarchical) tensor factorization as (deep) circuits highlights how circuit units can be naturally grouped together by type and scope into layers, as hinted already in Fig. 2. This perspective presents a new opportunity: defining and representing certain circuit structures as tensorized computational graphs. While circuits in the literature are defined in terms of scalar computational units, sum, product and inputs and single connections (Def. 2), many successful implementations of circuits nowadays already group units into tensors (Vergari et al., 2019a; Peharz et al., 2020c; a; Liu & Van den Broeck, 2021b; Loconte et al., 2024a) with the goal of speeding up computation by using the acceleration provided by GPUs. Following these ideas, we now provide a general tensorized circuit definition that offers a modular way to build overparameterized circuit architectures. This will allow us to design a single learning pipeline that subsumes many existing architectures (Section 4), and also suggest a way to create novel ones by mixing and reusing small “blocks”.
@@
Note that if a sum layer $\bm{\mathrm{\ell}}$ receives only one input vector, i.e., $|{\mathsf{in}}(\bm{\mathrm{\ell}})|=1$, then it simply computes $\bm{\mathrm{W}}\bm{\mathrm{\ell}}_{1}({\mathsf{sc}}(\bm{\mathrm{\ell}}_{1}))$. Furthermore, we retrieve the previous scalar unit-wise definition by setting $K$, the size of each layer, to 1. The above four types of layers constitute the basic “Lego blocks” that we will later use to create more sophisticated layers (Section 4.3, Section 5) and reproduce all modern circuit architectures (Table 1).
(a) input layer
(b) Hadamard layer
(c) Kronecker layer
(d) sum layer
@@ 2.4 Tensor Networks as Deep Circuits
Tensor networks (TNs) are often the preferred way to represent hierarchical tensor factorizations in fields such as physics and quantum computing (Markov & Shi, 2008; Schollwoeck, 2010). TNs come with a graphical language – Penrose notation – to encode tensor dot products in a compact graphical formalism (also called tensor contractions). See Orús (2013) for a review. Perhaps, the most popular TN factorization is the matrix-product state (MPS) (Pérez-García et al., 2007), also called tensor-train factorization (TT) (Oseledets, 2011; Glasser et al., 2019; Novikov et al., 2021). For instance, given a tensor $\bm{\mathcal{T}}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$, its rank-$R$ MPS/TT factorization is defined in element-wise notation as
@@
where $\bm{\mathrm{A}}^{(1)}\in\mathbb{R}^{I_{1}\times R}$, $\bm{\mathrm{A}}^{(d)}\in\mathbb{R}^{I_{d}\times R}$, and $\bm{\mathcal{A}}^{(j)}\in\mathbb{R}^{I_{j}\times R\times R}$ with $1<j<d$. That is, an MPS factorization decomposes $\bm{\mathcal{T}}$ into the complete contraction of a chain of smaller tensors $\bm{\mathrm{A}}^{(1)}$, $\bm{\mathrm{A}}^{(d)}$, and $\{\bm{\mathcal{A}}^{(j)}\}_{j=2}^{d-1}$. Fig. 7(a) shows an example of a MPS/TT represented in Penrose graphical notation, i.e., where nodes denote the tensors $\bm{\mathrm{A}}^{(1)},\bm{\mathcal{A}}^{(2)},\ldots,\bm{\mathcal{A}}^{(d-1)},\bm{\mathrm{A}}^{(d)}$, edges denote summations over shared indices, and $X_{1},\ldots,X_{d}$ denote the tensor indices whose assignment yield the corresponding tensor entry. Loconte et al. (2024a) showed how an MPS can be represented as deep tensorized circuits by encoding summations and products in Eq. 7 into sum and (Hadamard) product layers, respectively.
Figure 7: A MPS/TT represented as a deep tensorized circuit with Hadamard product layers (b). To obtain the parameters of the circuit, the tensor $\bm{\mathcal{A}}^{(2)}$ in the MPS/TT (a, showed in Penrose graphical notation) is firstly factorized into matrices $\bm{\mathrm{V}}^{(2)},\bm{\mathrm{B}},\bm{\mathrm{C}}$ through a CANDECOMP/PARAFAC decomposition (Carroll & Chang, 1970). Then, $\bm{\mathrm{V}}^{(1)},\bm{\mathrm{V}}^{(3)},\bm{\mathrm{W}}$ are obtained as in the figure (a). See Loconte et al. (2024a) for the detailed circuit construction. In (b) we denote with $\bm{1}$ a row-matrix whose entries are all ones.
Proposition 3 (MPS as deep tensorized circuits (Loconte et al., 2024a)).
@@
In Fig. 7 we show a tensorized circuit representing a MPS/TT over variables $\bm{\mathrm{X}}=\{X_{1},X_{2},X_{3}\}$, and, as detailed in the proof of Proposition 3 in Loconte et al. (2024a), the parameters of its input and dense layers are obtained by decomposing the tensors $\{\bm{\mathcal{A}}^{(j)}\}_{j=2}^{d-1}$ of the MPS/TT. Similarly to the tensorized circuit representation of hierarchical Tucker (2), 3 yields a tensorized circuit that is structured decomposable (Def. 6). Structured-decomposability is the crucial property in MPS/TTs that allows to perform certain operations over them tractably, for instance squaring them as to recover a Born machine – a probabilistic model devised to simulate quantum many-body systems in physics (Orús, 2013; Glasser et al., 2019). Understanding this enables practitioners to design alternative Born machine architectures that are not limited to a sequence of tensor operations as encoded in a “linear” RG, without having to prove the tractability of the square operation over these architectures from scratch (Shi et al., 2005). This is one of the opportunities we highlighted for hierarchical tensor factorizations once represented as circuits (Section 2.2 and Section 2.2). Further opportunities will be presented in the next section and directly translates to TNs as well as classical tensor factorizations.
@@
Until now, we discussed the generic case of decomposing a real-valued tensor. On the other hand, tensor factorizations that are tailored for non-negative data (e.g. images), called non-negative tensor factorizations, factorize tensors into non-negative factors that can be interpreted much more easily (Cichocki & Phan, 2009). Next, we connect non-negative tensor factorizations methods to the literature of circuits for probabilistic modeling, which allows us to interpret them as deep latent-variable models. In addition, by bridging non-negative tensor factorizations and their representation as (deep) circuits, we showcase a number of future research opportunities related to both parameterizing tensor factorizations and performing probabilistic inference with them.

### v2

@@ 2.2 Hierarchical Tensor Factorizations are Deep Circuits
Opportunity 1. A wider choice of factorization structures Refer to caption Figure 5: Region nodes can be shared between partitionings in a DAG RG. Def. 4 allows for arbitrary DAGs and arbitrarily scoped-regions, while hierarchical tensor factorizations are usually presented in terms of RGs with a tree structure having region leaves containing exactly one variable (e.g., the RG in Fig. 3), which are sometimes called dimension trees or mode cluster trees (Grasedyck, 2010) in the tensor factorization community, and often vtree in circuit literature (Pipatsrisawat & Darwiche, 2008; Kisa et al., 2014; Wedenig et al., 2024a). More intricate RG structures can increase the expressiveness as well as flexibility in building deep circuits/hierarchical factorizations. Intuitively, we can share factor matrices among multiple factorizations, and therefore reduce the number of model parameters, making a more space-efficient implementation possible. See Peharz et al. (2020c; a) for more details. We provide an example of such a RG in Fig. 10 and Fig. 5 here on the left. The reader can check that this RG encodes the hierarchical scope partitioning of the decomposable circuit in figure Fig. 1. Fig. 9 then illustrates a fragment of this RG and shows how tensor factorizations conforming to it can be constructed as circuits in Section 4. The circuit literature provides several ways to build RGs that are suitable for certain data modalities (e.g., image, sequence, tabular data), which can also be learned from data. Section 4.1 provides an overview of such techniques.
Figure 6: Tensorized circuits can encode novel hierarchical multilinear factorizations by mixing different structures and layers. Section 2.3 and Fig. 7 formalize and illustrate our tensorized circuit formalism, respectively. The figure on the left shows a tensorized circuit over variables $\bm{\mathrm{X}}=\{X_{1},X_{2},X_{3}\}$ encoding an allowed multilinear factorization of a three-dimensional tensor (as it is smooth and decomposable, see Def. 8). Note that each input layer has its own factor matrix $\bm{\mathrm{V}}^{(j)}$ with $j\in[6]$, and the architecture consists of a mix of Hadamard, Kronecker and sum inner layers. Overall, this tensorized circuit do not map to a known hierarchical factorization. The figure on the right shows a similar tensorized circuit, where the factor matrices $\bm{\mathrm{V}}^{(2)}$ and $\bm{\mathrm{V}}^{(3)}$ are instead shared while still encoding an allowed multilinear factorization.
@@
Imposing a particular factorization structure by leveraging a RG, and picking a particular parameterization for each region in it (as it will be discussed in Section 4), represents one way to encode novel hierarchical factorizations that do not correspond to existing ones. Fig. 6 shows some examples. There, we represent circuits in a layer-wise formalism as described later in Section 2.3. Note that instantiating tensor factorizations from RGs defined as above preserve decomposability, and that circuits built from RGs in the literature are typically also smooth (Def. 3). Hierarchical Tucker and its variants are also smooth and decomposable and therefore support the tractable computation of a number of (probabilistic) inference tasks (Section 3). These hierarchical factorizations (and the corresponding deep circuits) that follow a tree-shaped RG with univariate leaves satisfy an additional structural property, called structured-decomposability. Structured decomposability enables the tractable computation of harder operations for which smoothness and decomposability are not enough. For instance, squaring particular tensor factorizations formalized in the graphical language of tensor networks, known as the Born rule in physics (Feynman, 1987; Glasser et al., 2019) (see also Section 2.4). We define structured decomposability below.
@@
We can easily check that hierarchical Tucker yields a structured decomposable circuit, as it is obtained by stacking Tucker factorizations (which are computed by decomposable circuits) based on a tree RG, which in turn synchronizes all product units to decompose in the same way. We emphasize that eliciting the few structural properties that can explain the tractable computation of many different quantities of interest can help save effort aimed at (re)discovering and (re)engineering algorithms for specific hierarchical factorizations.
@@
Opportunity 2. Efficient Compositional Operations over Factorizations Given one or more tensor factorizations appearing as operands in a computation of interest, how can we automatically devise a tractable algorithm for it without having to materialize the exponentially large tensor operands? The circuit literature holds the answer and offers other structural properties that can unlock the tractable computation of many complex inference scenarios, in a reusable fashion. E.g., when two deep circuits conform to the same tree RG, they are said to be compatible (Vergari et al., 2021). Given two compatible hierarchical tensor factorizations $p$ and $q$ over $\bm{\mathrm{X}}$, one can compute general expectations of the form $\sum\nolimits_{\bm{\mathrm{x}}}p(\bm{\mathrm{x}})q(\bm{\mathrm{x}})$ (expectations) in closed form in time $\mathcal{O}(|p||q|)$, where $|p|$ and $|q|$ are the size of the corresponding circuits encoding such factorizations. On the other hand, maximization problems as in maximum-a-posteriori inference $\max\nolimits_{\bm{\mathrm{y}}}p(\bm{\mathrm{y}},\bm{\mathrm{E}}=\bm{\mathrm{e}})$ (MAP inference) where $\bm{\mathrm{e}}$ is the evidence assignment to variables $\bm{\mathrm{E}}\subset\bm{\mathrm{X}}$, and $\bm{\mathrm{y}}$ is the assignment to the remaining variables $\bm{\mathrm{Y}}=\bm{\mathrm{X}}\setminus\bm{\mathrm{E}}$ for which we want to maximize $p$, can be solved exactly and efficiently if $p$ is a decomposable circuit that supports an additional property, determinism (Darwiche, 2009). In a nutshell, sum units in a deterministic circuit receive inputs from functions with disjoint support (see Choi et al. (2020) for details). While determinism is a consolidated property in the circuit literature, it is off the radar for (hierarchical) tensor factorizations. Furthermore, the circuit literature provides a systematic way to quickly devise the tractability conditions for a given mathematical expression that involves sums, products, powers, exponentials and logarithms, and therefore automatically distill corresponding tractable algorithms (Vergari et al., 2021). For example, if one wants to compute Rényi’s $\alpha$-divergence between two factorizations $p$ and $q$ over variables $\bm{\mathrm{X}}$, for $\alpha\in\mathbb{N}$, defined as $(1-\alpha)^{-1}\log\sum\nolimits_{\bm{\mathrm{x}}}p^{\alpha}(\bm{\mathrm{x}})q^{1-\alpha}(\bm{\mathrm{x}}),\$ ($\alpha$-divergence) then this can be done quickly if $p$ and $q$ can be represented as smooth, decomposable and compatible circuits and $q$ is also deterministic. Vergari et al. (2021) show how to automatically distill the tractable computation of more information-theoretic quantities.
@@ 2.3 Representing Circuits in a Tensorized Formalism
(a) input layer
(b) Hadamard layer
(c) Kronecker layer
(d) sum layer
Representing (hierarchical) tensor factorization as (deep) circuits highlights how circuit units can be naturally grouped together by type and scope into layers, as hinted already in Fig. 2. This perspective presents a new opportunity: defining and representing certain circuit structures as tensorized computational graphs. While circuits in the literature are defined in terms of scalar computational units, sum, product and inputs and single connections (Def. 2), many successful implementations of circuits nowadays already group units into tensors (Vergari et al., 2019a; Peharz et al., 2020c; a; Liu & Van den Broeck, 2021b; Loconte et al., 2024) with the goal of speeding up computation by using the acceleration provided by GPUs. Following these ideas, we now provide a general tensorized circuit definition that offers a modular way to build overparameterized circuit architectures. This will allow us to design a single learning pipeline that subsumes many existing architectures (Section 4), and also suggest a way to create novel ones by mixing and reusing small “blocks”.
@@
Note that if a sum layer $\bm{\mathrm{\ell}}$ receives only one input vector, i.e., $|{\mathsf{in}}(\bm{\mathrm{\ell}})|=1$, then it simply computes $\bm{\mathrm{W}}\bm{\mathrm{\ell}}_{1}({\mathsf{sc}}(\bm{\mathrm{\ell}}_{1}))$. Fig. 7 illustrates the layer types of a tensorized circuit, together with the unit-wise representation (Def. 2). Furthermore, we retrieve the previous scalar unit-wise definition by setting $K$, the size of each layer, to 1. The above four types of layers constitute the basic “Lego blocks” that we will later use to create more sophisticated layers (Section 4.3, Section 5) and reproduce all modern circuit architectures (Table 1).
@@
Figure 8: A MPS/TT represented as a deep tensorized circuit with Hadamard product layers (b). To obtain the parameters of the circuit, the tensor $\bm{\mathcal{A}}^{(2)}$ in the MPS/TT (a, showed in Penrose graphical notation) is firstly factorized into matrices $\bm{\mathrm{V}}^{(2)},\bm{\mathrm{B}},\bm{\mathrm{C}}$ through a CANDECOMP/PARAFAC decomposition (Carroll & Chang, 1970). Then, $\bm{\mathrm{V}}^{(1)},\bm{\mathrm{V}}^{(3)},\bm{\mathrm{W}}$ are obtained as in the figure (a). See Loconte et al. (2024) for the detailed circuit construction. In (b) we denote with $\bm{1}$ a row-matrix whose entries are all ones.
@@ 2.4 Tensor Networks as Deep Circuits
Tensor networks (TNs) are often the preferred way to represent hierarchical tensor factorizations in fields such as physics and quantum computing (Markov & Shi, 2008; Schollwoeck, 2010; Biamonte & Bergholm, 2017). TNs come with a graphical language – Penrose notation – to encode tensor dot products in a compact graphical formalism (also called tensor contractions). See Orús (2013) for a review. Perhaps, the most popular TN factorization is the matrix-product state (MPS) (Pérez-García et al., 2007), also called tensor-train factorization (TT) (Oseledets, 2011; Glasser et al., 2019; Novikov et al., 2021). For instance, given a tensor $\bm{\mathcal{T}}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$, its rank-$R$ MPS/TT factorization is defined in element-wise notation as
@@
where $\bm{\mathrm{A}}^{(1)}\in\mathbb{R}^{I_{1}\times R}$, $\bm{\mathrm{A}}^{(d)}\in\mathbb{R}^{I_{d}\times R}$, and $\bm{\mathcal{A}}^{(j)}\in\mathbb{R}^{I_{j}\times R\times R}$ with $1<j<d$. That is, an MPS factorization decomposes $\bm{\mathcal{T}}$ into the complete contraction of a chain of smaller tensors $\bm{\mathrm{A}}^{(1)}$, $\bm{\mathrm{A}}^{(d)}$, and $\{\bm{\mathcal{A}}^{(j)}\}_{j=2}^{d-1}$. Fig. 8(a) shows an example of a MPS/TT represented in Penrose graphical notation, i.e., where nodes denote the tensors $\bm{\mathrm{A}}^{(1)},\bm{\mathcal{A}}^{(2)},\ldots,\bm{\mathcal{A}}^{(d-1)},\bm{\mathrm{A}}^{(d)}$, edges denote summations over shared indices, and $X_{1},\ldots,X_{d}$ denote the tensor indices whose assignment yield the corresponding tensor entry. Loconte et al. (2024) showed how an MPS can be represented as a deep tensorized circuit by encoding summations and products in Eq. 7 into sum and (Hadamard) product layers, respectively.
Proposition 3 (MPS as deep tensorized circuits (Loconte et al., 2024)).
@@
In Fig. 8 we show a tensorized circuit representing a MPS/TT over variables $\bm{\mathrm{X}}=\{X_{1},X_{2},X_{3}\}$, and, as detailed in the proof of Proposition 3 in Loconte et al. (2024), the parameters of its input and dense layers are obtained by decomposing the tensors $\{\bm{\mathcal{A}}^{(j)}\}_{j=2}^{d-1}$ of the MPS/TT. Similarly to the tensorized circuit representation of hierarchical Tucker (2), 3 yields a tensorized circuit that is structured decomposable (Def. 6). Structured-decomposability is the crucial property in MPS/TTs that allows to perform certain operations over them tractably, for instance squaring them as to recover a Born machine – a probabilistic model devised to simulate quantum many-body systems in physics (Orús, 2013; Glasser et al., 2019). Understanding this enables practitioners to design alternative Born machine architectures that are not limited to a sequence of tensor operations as encoded in a “linear” RG, without having to prove the tractability of the square operation over these architectures from scratch (Shi et al., 2005). This is one of the opportunities we highlighted for hierarchical tensor factorizations once represented as circuits (Section 2.2 and Section 2.2). Further opportunities will be presented in the next section and directly translates to TNs as well as classical tensor factorizations.
@@
Until now, we discussed the generic decomposition of a real-valued tensor. However, tensor factorizations that are tailored for non-negative data (e.g. images), called non-negative tensor factorizations, factorize tensors into non-negative factors that can be easily interpreted (Cichocki & Phan, 2009). In Section 3, we connect non-negative tensor factorizations to the literature of circuits for probabilistic modeling, which allows us to interpret them as deep latent-variable models. In addition, by bridging non-negative tensor factorizations and their representation as (deep) circuits, we showcase future research opportunities related to both parameterizing tensor factorizations and performing probabilistic inference with them.